The total revenues of consumer electronics since 2000 are approximated by the function f(t) = 78.5 * 1.1(12 3), where t = 3 corresponds to 2003 and f(t) is in billions of dollars. Find the revenue in the given years below. (a) 2006 (b) 2008 (c) If the function remains accurate, what will the revenues be in 2013? (a) In 2006, the total revenues were approximately S bilion. (Round to two decimal places as needed.)
Answer #1

f(t)=78.5*1.1^{t}

(a) in 2006 the total revenues were =78.5*1.1^{6}
=139.07 billion

(b) in 2008 the total revenues were =78.5*1.1^{8}
=168.27 billion

(c) in 2013 the total revenues were =78.5*1.1^{13}
=271.00 billion